# Select, then Balance: A Plug-and-Play Framework for Exogenous-Aware Spatio-Temporal Forecasting

### v1

@@ 4.6. Efficiency & Lightweight Study (RQ5)
We compared the differences in training time, memory footprint, and performance between ExoST and other time series or spatio-temporal forecasting models that incorporate exogenous variables on the AQI-19 dataset. Due to varying convergence epochs, for a fair comparison, we only present the average training time for the first 100 epochs. As shown in Figure 7, compared to several other methods, ExoST demonstrates a balance of performance, efficiency, and model parameter count. Furthermore, we can observe that: ❶ compared to time series models with different backbones, despite its lightweight model design, it leads to poor performance. ❷ compared to spatio-temporal method and the general exogenous variables modeling method with the same backbone (i.e., MAGCRN, ChronosX), we achieved further performance improvement while maintaining similar parameters and training time.
5. Conclusion and Future Work
@@
This paper investigates the challenges of modeling exogenous variables for spatio-temporal forecasting and explores effective implementation options. We propose a novel framework, ExoST, which follows a “select, then balance” paradigm. It first dynamically selects and restructures salient signals via a latent space gated expert module to address inconsistent variable effects. It then incorporates a dual-branch architecture with a context-aware weighting mechanism to address imbalanced type effects. Extensive experiments demonstrate its effectiveness, universality, and robustness. In future work, we aim to explore how to broaden the modal availability of exogenous variables and their feasibility in data-scarce scenarios.
Supplementary material
Select, then Balance: A Plug-and-Play Framework for Exogenous-Aware Spatio-Temporal Forecasting
[…]
[sections] Table of Contents
[sections]l1
Appendix A Experimental Details
A.1. Datasets Details
@@
The statistics of the two real-world datasets used in this paper are summarized in Table 1.
| Attribute | Madrid-19 | Madrid-22 | ||||
| Time range | 01/01/2019 – 30/06/2019 | 01/01/2022 – 30/06/2022 | ||||
| Frames | 4344 | 4344 | ||||
| Sampling Rate | 1 hour | 1 hour | ||||
| Data points | 2.03 M | 2.03 M | ||||
| Variables | 20 | 20 | ||||
| Variable Domain |
@@ A.2. Baseline Details
A.2.1. Exogenous Variable Modeling Methods
@@
TiDE (Das et al., 2023): TiDE is an encoder–decoder model built on multilayer perceptrons (MLPs) that fuses historical time-series data and covariates via dense MLP residual blocks to map them into low-dimensional representations, then generates forecast sequences by incorporating future covariates. https://github.com/google-research/google-research/tree/master/tide
TimeXer (Wang et al., 2024): TimeXer is a Transformer model tailored for time series forecasting with exogenous variables: it captures temporal dependencies in the endogenous series via patch-level self-attention and effectively integrates exogenous information through variable-level cross attention and learnable global endogenous tokens to produce high-quality predictions of the target series. https://github.com/thuml/TimeXer
NBEATSx (Olivares et al., 2023): NBEATSx extends the original N-BEATS by incorporating convolutional encoding substructures for time-varying and static exogenous variables within a hierarchical stacked residual block architecture, enabling interpretable decomposition of trend, seasonality, and exogenous factors for high-quality target-series forecasting. https://github.com/cchallu/nbeatsx
@@
MAGCRN (Giganti et al., 2024): MAGCRN is a spatio-temporal model that extends AGCRN into a dual-path architecture, using separate AGCRN blocks to encode historical observations and future exogenous covariates. https://github.com/polimi-ispl/MAGCRN
@@
ChronosX (Arango et al., 2025): ChronosX extends MAGCRN by integrating two exogenous adaptation modules: Instance-wise Interpolation Block (IIB) and Offset Integration Block (OIB). IIB adaptively interpolates exogenous inputs across time steps, while OIB injects time-dependent bias offsets into the output layer, enabling more flexible integration of future covariates. Due to the limited public availability of the code, we implemented the IIB and OIB modules based on the formulas provided in the referenced paper. https://github.com/amazon-science/chronos-forecasting/tree/chronosx
A.2.2. Spatio-Temporal Backbone Models
@@
AGCRN (Bai et al., 2020): AGCRN is an adaptive graph-based recurrent model that captures fine-grained spatio-temporal dependencies by jointly learning node-specific patterns and data-driven graph structures, eliminating the need for pre-defined adjacency matrices.
GWNet (Wu et al., 2019): GWNet is a spatial-temporal model that combines diffusion-based graph convolutions with dilated causal temporal convolutions, and incorporates a learnable self-adaptive adjacency matrix to capture latent spatial relationships from data.
GGNN (Satorras et al., 2022): GGNN is a recurrent graph-based model that applies gated update mechanisms to iteratively refine node representations through message passing, enabling effective modeling of long-range structural dependencies.
GRUGCN (Gao and Ribeiro, 2022): GRUGCN is a spatio-temporal neural architecture that integrates graph convolutions for spatial feature extraction and gated recurrent units for sequential modeling, effectively capturing dynamic dependencies in traffic forecasting scenarios.
STGCN (Yu et al., 2017): STGCN is a graph-based temporal forecasting model that integrates graph convolutional layers with temporal convolution modules to jointly learn spatial and temporal dependencies in structured time series data.
DCRNN (Li et al., 2017): DCRNN is a graph-based recurrent forecasting model that captures spatial dependencies via diffusion convolutions on directed graphs and models temporal dynamics through a sequence-to-sequence recurrent architecture.
A.3. Protocol Details
A.3.1. Metrics Detail.
@@ A.3. Protocol Details
Appendix B More Results
B.1. More Result on Universality Study
@@
[…]
@@
[…]
We provide more experimental results on universality (i.e., for 1-day and 2-day scenarios), as shown in Tables 2 and 3. We observe that while the ExoST exhibits some negative gains in the 1-day scenario, it exhibits improved performance as the prediction length increases, demonstrating the superior generalization of our method. The remaining observations show similar analytical results to those in the main text.
B.2. More Result on Robustness Study
@@
| Zero | 20%-Mask | ||||||||||||
| 40%-Mask | |||||||||||||
| 60%-Mask | |||||||||||||
| 80%-Mask | |||||||||||||
| Random | 20%-Mask | ||||||||||||
| 40%-Mask | |||||||||||||
| 60%-Mask | |||||||||||||
| 80%-Mask | |||||||||||||
| No Masking | |||||||||||||
@@
| Zero | 20%-Mask | ||||||||||||
| 40%-Mask | |||||||||||||
| 60%-Mask | |||||||||||||
| 80%-Mask | |||||||||||||
| Random | 20%-Mask | ||||||||||||
| 40%-Mask | |||||||||||||
| 60%-Mask | |||||||||||||
| 80%-Mask | |||||||||||||
| No Masking | |||||||||||||
B.3. More Result on Mechanism Study
B.3.1. Ablation Study.
@@ B.3. More Result on Mechanism Study
[…]
From the data perspective, we also evaluated the impact of exogenous variables on ExoST’s performance for 2-day and 3-day forecasting on the AQI-19 , as shown in Tables 6 and Tables 7. We have further observation: ❶ Consistent with the 1-day findings, future covariates outperform past covariates, while data variables alone produce limited accuracy but enhance performance when combined with others, with the best results achieved by integrating all variables, demonstrating ExoST’s ability to synthesize diverse temporal information. ❸ In 2-day forecasting, combining past and future covariates significantly improves percentage-based metrics, indicating a balanced contribution from historical and forward-looking signals. ❷ For 3-day forecasting, combining future and data covariates almost matches the full model, highlighting the significant role of the immediate temporal context in short-term predictions. ❹ Across 1-day to 3-day horizons, the most critical trend is increasing the importance of the past covariates as the forecast horizon extends, complementing the dominant predictive power of future covariates.
B.3.2. Strategy Study.
@@
Simple Weight: Each stream is processed by its own encoder 3.2. The final prediction employs fixed, equal weights:
@@
Attention: Let denote the intermediate hidden states produced by the two branches before the final read-out layer. For each directional attention we use shared projection matrices .
Future Past
@@
As Figures 10 and 11 show, we can observe that: ❶ In the 2-day horizon, ExoST maintains its advantage, while Simple Weight and Learnable Weight exhibit amplified weaknesses: the former’s static weighting fails to handle increased uncertainty in longer forecasts, and the latter, despite its adaptable weights, amplifies noise from future exogenous signals due to lacking day-to-day context modeling. ❷ Althougn all methods show increasing errors across horizons of 1-day to 3-day, ExoST consistently suppresses error accumulation better than alternatives.
Appendix C More Disscuison
C.1. Limitation

### v2

@@ 5.6 Efficiency & Lightweight Study (RQ5)
We evaluated the training time, memory footprint, and performance of ExoST against other forecasting models incorporating exogenous variables on the AQI-19 dataset. To ensure fair comparison given varying convergence rates, we report the average training time (first 100 epochs). As shown in Figure 9, ExoST strikes a favorable balance between performance, efficiency, and parameter count. Specifically: ❶ Compared to time series models, purely lightweight designs often lead to inferior performance. ❷ Compared to ST and general exogenous variable modeling methods using the same backbone (i.e., MAGCRN, ChronosX), our approach achieves superior performance while maintaining comparable parameter counts and training times.
6 Conclusion and Future Work
@@
This paper presents the first systematic study of the challenges inherent in modeling exogenous variables for spatio-temporal forecasting and explores effective implementation strategies. We propose a simple yet effective general framework, ExoST. This framework follows a “select, then balance” paradigm, which effectively addresses the problems of inconsistent variable effects and unbalanced type effects. Extensive experiments demonstrate its effectiveness, universality, and robustness. In future work, we aim to explore how to expand the modality availability of exogenous variables and their feasibility in extremely data-scarce scenarios.
Acknowledgment
This work is mainly supported by the National Natural Science Foundation of China (No. 62402414). This work is also supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2024A03J0620), Guangzhou Municipal Science and Technology Project (No. 2023A03J0011), the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628), and a grant from State Key Laboratory of Resources and Environmental Information System, and Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No. 2023B1212010007).
Appendix
Appendix A More Analysis
A.1 More Analysis on Select Module
We provide theoretical insights into why the proposed select stage (conditional embedding + latent-space gated experts) improves exogenous-aware spatio-temporal forecasting, addressing the challenge of inconsistent variable effects. Our analysis is backbone-agnostic and focuses on representation, approximation, and optimization perspectives.
✦Input-dependent linear operator and regime-adaptive processing. The gated expert selector in Eq. (4) can be reformulated as an input-dependent linear operator:
where is a simplex-valued gate. Unlike a static projection that enforces a global transformation, Eq. (A.1) implements a dynamic filter whose parameters evolve with local spatio-temporal contexts encoded in . This explicitly models the inconsistent effects, where the contribution of endogenous channels varies across different weather conditions or temporal regimes.
✦Latent regime approximation. We postulate that the data-generating process involves a continuous latent regime variable (e.g., a continuous spectrum from clear to extreme weather). The predictive distribution can be viewed as an integral over these regimes:
Our module approximates this integral via discretization. The conditional embedding aligns heterogeneous sources, enabling the gate to approximate the regime posterior . The experts then serve as support points in the function space, effectively decomposing the complex multimodal mapping into a mixture of local experts:
✦Reduced gradient interference via soft routing. Inconsistent effects also manifest as gradient conflicts: samples from different contexts induce updates in opposing directions when sharing parameters. For an expert parameter (e.g., ), the per-sample gradient satisfies
where is the gate weight for sample . Thus, samples primarily update the experts they activate, which decomposes conflicting gradients across experts and reduces negative transfer. Empirically, this is consistent with smoother gate trajectories and clearer expert specialization (Fig. 4), improving optimization stability and downstream forecasting accuracy.
Overall, the key advantage of the Select Module is that it explicitly incorporates exogenous-condition-induced variations in variable effects into both the representation learning and routing mechanisms. Specifically, the conditional embedding first fuses endogenous histories with heterogeneous, multi-source exogenous signals into context-dependent latent representations, thereby establishing a unified semantic basis for subsequent selection. Building on this latent space, the gated experts further instantiate an input-dependent dynamic operator, enabling the model to automatically activate more appropriate processing pathways across different time steps, spatial nodes, and exogenous contexts. Meanwhile, soft routing assigns the learning signal from each contextual sample primarily to its corresponding expert parameters, effectively alleviating gradient interference and negative transfer that commonly arise under shared-parameter training, which in turn promotes context-specialized expert division of labor and improves optimization efficiency. Consequently, by combining conditional representation with expert selection, the Select stage realizes adaptive modeling of inconsistent variable effects and provides the downstream backbone with more robust and controllable input features.
A.2 More Analysis on Balance Module
Although existing methods (wang2024timexer; giganti2024back; Arango2025ChronosX) have explored various strategies for balancing past and future exogenous variables, our approach introduces a fundamental distinction in its design philosophy. We theoretically scrutinize the discrepancies among these methods through the lens of functional complexity and the representational capacity of their respective function spaces:
Fixed Weighting Strategy: This approach operates as a rudimentary static linear mapping, formalized as:
The representational capacity of this method is severely constrained, as the dimension of the function space is , governed exclusively by the scalar hyperparameter . It imposes a rigid interpolation that is agnostic to data distribution.
Learnable Weighting Strategy: While this method introduces optimization, it fundamentally remains a linear mapping within the family of affine transformations:
Although the parameters are learnable, the objective is to converge upon a single, global optimal weighted balance across the entire training set. Consequently, the learned weights remain static during the inference phase and cannot adapt to the specific variations of individual samples.
Context-Aware Balancing (Ours): In stark contrast, our proposed mechanism functions as an input-dependent, non-linear mapping:
Theoretically, this formulation is capable of approximating arbitrary fusion functions. By conditioning the weights on the input itself, the model achieves context sensitivity at inference time, allowing for the dynamic modulation of the fusion strategy based on immediate data characteristics.
Therefore, our methodology diverges significantly from prior static strategies by offering superior dynamic adaptability. To validate this theoretical advantage empirically, Section 5.5 presents a comprehensive ablation study comparing our approach against degenerate fixed-weight baselines, learnable-weight strategies, and attention-based mechanisms. The empirical results robustly demonstrate the efficacy and necessity of our context-aware balancing strategy.
Appendix B Experimental Details
B.1 Datasets Details
@@
The statistics of the two real-world datasets used in this paper are summarized in Table B.1.
| Attribute | Madrid-19 | Madrid-22 |
| Time range | 01/01/2019 – 30/06/2019 | 01/01/2022 – 30/06/2022 |
| Frames | 4344 | 4344 |
| Sampling Rate | 1 hour | 1 hour |
| Data points | 2.03 M | 2.03 M |
| Variables | 20 | 20 |
| Variable Domain | Air quality, Traffic, Meteorological, Date | Air quality, Traffic, Meteorological, Date |
@@ B.2 Baseline Details
B.2.1 Exogenous Variable Modeling Methods
@@
TiDE (das2023long): TiDE is an encoder–decoder model built on multilayer perceptrons (MLPs) that fuses historical time-series data and covariates via dense MLP residual blocks to map them into low-dimensional representations, then generates forecast sequences by incorporating future covariates. https://github.com/google-research/google-research/tree/master/tide
TimeXer (wang2024timexer): TimeXer is a Transformer model tailored for time series forecasting with exogenous variables: it captures temporal dependencies in the endogenous series via patch-level self-attention and effectively integrates exogenous information through variable-level cross attention and learnable global endogenous tokens to produce high-quality predictions of the target series. https://github.com/thuml/TimeXer
NBEATSx (olivares2023neural): NBEATSx extends the original N-BEATS by incorporating convolutional encoding substructures for time-varying and static exogenous variables within a hierarchical stacked residual block architecture, enabling interpretable decomposition of trend, seasonality, and exogenous factors for high-quality target-series forecasting. https://github.com/cchallu/nbeatsx
CrossLinear (zhou2025crosslinear):CrossLinear is a Linear-based model that introduces a plug-and-play cross-correlation embedding module to capture time-invariant dependencies between endogenous and exogenous variables, utilizing patch-wise processing and a global linear head to effectively model both short-term and long-term temporal patterns for precise forecasting. https://github.com/mumiao2000/CrossLinear
DAG (qiu2025dagdualcausalnetwork): DAG is a general framework that utilizes a Dual Causal Network along temporal and channel dimensions, it explicitly models how historical exogenous variables affect future exogenous variables (temporal) and historical endogenous variables (channel), and then injects these discovered causal relationships to improve the prediction of future endogenous variables. https://github.com/decisionintelligence/DAG
@@
MAGCRN (giganti2024back): MAGCRN is a spatio-temporal model that extends AGCRN into a dual-path architecture, using separate AGCRN blocks to encode historical observations and future exogenous covariates. https://github.com/polimi-ispl/MAGCRN
@@
ChronosX (Arango2025ChronosX): ChronosX extends MAGCRN by integrating two exogenous adaptation modules: Instance-wise Interpolation Block (IIB) and Offset Integration Block (OIB). IIB adaptively interpolates exogenous inputs across time steps, while OIB injects time-dependent bias offsets into the output layer, enabling more flexible integration of future covariates. Due to the limited public availability of the code, we implemented the IIB and OIB modules based on the formulas provided in the referenced paper. https://github.com/amazon-science/chronos-forecasting/tree/chronosx
B.2.2 Spatio-Temporal Backbone Models
@@
AGCRN (Bai2020Adaptive): AGCRN is an adaptive graph-based recurrent model that captures fine-grained spatio-temporal dependencies by jointly learning node-specific patterns and data-driven graph structures, eliminating the need for pre-defined adjacency matrices.
GWNet (wu2019graph): GWNet is a spatial-temporal model that combines diffusion-based graph convolutions with dilated causal temporal convolutions, and incorporates a learnable self-adaptive adjacency matrix to capture latent spatial relationships from data.
GGNN (Satorras2022Multivariate): GGNN is a recurrent graph-based model that applies gated update mechanisms to iteratively refine node representations through message passing, enabling effective modeling of long-range structural dependencies.
GRUGCN (Gao2022On): GRUGCN is a spatio-temporal neural architecture that integrates graph convolutions for spatial feature extraction and gated recurrent units for sequential modeling, effectively capturing dynamic dependencies in traffic forecasting scenarios.
STGCN (yu2017spatio): STGCN is a graph-based temporal forecasting model that integrates graph convolutional layers with temporal convolution modules to jointly learn spatial and temporal dependencies in structured time series data.
DCRNN (li2017diffusion): DCRNN is a graph-based recurrent forecasting model that captures spatial dependencies via diffusion convolutions on directed graphs and models temporal dynamics through a sequence-to-sequence recurrent architecture.
B.3 Protocol Details
B.3.1 Metrics Detail.
@@ B.3 Protocol Details
Appendix C More Results
C.1 More Result on Universality Study
@@
| AQI-19 | MAE | 10.41±0.12 | |||||||||||||||||
| RMSE | 15.51±0.16 | ||||||||||||||||||
| MAPE(%) | 44.65±0.08 | ||||||||||||||||||
| MRE(%) | 40.34±0.47 | ||||||||||||||||||
| Speed-19 | MAE | 3.80±0.01 | |||||||||||||||||
| RMSE | 31.85±0.02 | ||||||||||||||||||
| MAPE(%) | 9.31±0.03 | ||||||||||||||||||
| MRE(%) | 9.90±0.03 | ||||||||||||||||||
| AQI-22 | MAE | 7.94±0.29 | |||||||||||||||||
| RMSE | 11.79±0.51 | ||||||||||||||||||
| MAPE(%) | 50.27±1.18 | ||||||||||||||||||
| MRE(%) | 41.18±1.83 | ||||||||||||||||||
| Intensity-22 | MAE | 61.87±3.80 | |||||||||||||||||
| RMSE | 105.44±5.53 | ||||||||||||||||||
| MAPE(%) | 31.74±1.69 | ||||||||||||||||||
| MRE(%) | 22.35±1.37 | ||||||||||||||||||
@@
| AQI-19 | MAE | 10.04±0.19 | |||||||||||||||||
| RMSE | 15.12±0.26 | ||||||||||||||||||
| MAPE(%) | 42.62±0.52 | ||||||||||||||||||
| MRE(%) | 38.84±0.75 | ||||||||||||||||||
| Speed-19 | MAE | 3.77±0.06 | |||||||||||||||||
| RMSE | 31.38±0.02 | ||||||||||||||||||
| MAPE(%) | 9.05±0.23 | ||||||||||||||||||
| MRE(%) | 9.83±0.15 | ||||||||||||||||||
| AQI-22 | MAE | 7.35±0.21 | |||||||||||||||||
| RMSE | 11.47±0.39 | ||||||||||||||||||
| MAPE(%) | 46.59±0.64 | ||||||||||||||||||
| MRE(%) | 39.35±1.11 | ||||||||||||||||||
| Intensity-22 | MAE | 50.53±3.94 | |||||||||||||||||
| RMSE | 89.10±7.36 | ||||||||||||||||||
| MAPE(%) | 26.11±2.67 | ||||||||||||||||||
| MRE(%) | 18.29±1.43 | ||||||||||||||||||
| Model | AGCRN | GWNet | GGNN | GRUGCN | STGCN | DCRNN | |||||||||||||
| w/ ExoST | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | |
| AQI-19 | MAE | 9.33±0.13 | |||||||||||||||||
| RMSE | 14.24±0.18 | ||||||||||||||||||
| MAPE(%) | 38.57±0.32 | ||||||||||||||||||
| MRE(%) | 35.85±0.71 | ||||||||||||||||||
| Speed-19 | MAE | 3.70±0.06 | |||||||||||||||||
| RMSE | 30.92±0.04 | ||||||||||||||||||
| MAPE(%) | 8.59±0.21 | ||||||||||||||||||
| MRE(%) | 9.64±0.16 | ||||||||||||||||||
| AQI-22 | MAE | 6.81±0.30 | |||||||||||||||||
| RMSE | 10.72±0.58 | ||||||||||||||||||
| MAPE(%) | 31.80±0.70 | ||||||||||||||||||
| MRE(%) | 36.44±1.62 | ||||||||||||||||||
| Intensity-22 | MAE | 29.74±1.21 | |||||||||||||||||
| RMSE | 48.88±1.95 | ||||||||||||||||||
| MAPE(%) | 17.63±0.97 | ||||||||||||||||||
| MRE(%) | 10.78±0.44 | ||||||||||||||||||
We provide more experimental results on universality (i.e., for 1-day and 2-day scenarios), as shown in Tables C.2 and C.3. We observe that while the ExoST exhibits some negative gains in the 1-day scenario, it exhibits improved performance as the prediction length increases, demonstrating the superior generalization of our method. The remaining observations show similar analytical results to those in the main text.
C.2 More Result on Robustness Study
@@
| Zero | 20% | ||||||||||||
| 40% | 9.23±0.01 | 9.84±0.01 | |||||||||||
| 60% | |||||||||||||
| 80% | |||||||||||||
| Random | 20% | ||||||||||||
| 40% | 31.40±1.10 | 54.03±1.13 | 11.38±0.40 | ||||||||||
| 60% | |||||||||||||
| 80% | 3.78±0.01 | 31.72±0.03 | 9.23±0.01 | 9.84±0.02 | |||||||||
| No Masking | 9.33±0.13 | 14.24±0.18 | 38.57±0.32 | 36.09±0.49 | 17.630.97 | ||||||||
@@
| Zero | 20% | 50.06±6.65 | 89.00±13.90 | 26.06±2.03 | 18.12±2.40 | ||||||||
| 40% | |||||||||||||
| 60% | |||||||||||||
| 80% | |||||||||||||
| Random | 20% | 9.32±0.05 | |||||||||||
| 40% | |||||||||||||
| 60% | |||||||||||||
| 80% | 9.32±0.04 | ||||||||||||
| No Masking | 10.04±0.19 | 15.12±0.26 | 42.62±0.52 | 38.84±0.75 | 3.80±0.01 | 31.78±0.02 | 9.32±0.03 | 9.89±0.03 | |||||
C.3 More Result on Mechanism Study
C.3.1 Ablation Study.
@@ C.3 More Result on Mechanism Study
| ✓ | ✓ | ✓ | 10.41±0.12 | 15.51±0.16 | 40.34±0.47 |
From the data perspective, we also evaluated the impact of exogenous variables on ExoST’s performance for 2-day and 3-day forecasting on the AQI-19 , as shown in Tables C.7 and Tables C.7. We have further observation: ❶ Consistent with the 1-day findings, future covariates outperform past covariates, while data variables alone produce limited accuracy but enhance performance when combined with others, with the best results achieved by integrating all variables, demonstrating ExoST’s ability to synthesize diverse temporal information. ❸ In 2-day forecasting, combining past and future covariates significantly improves percentage-based metrics, indicating a balanced contribution from historical and forward-looking signals. ❷ For 3-day forecasting, combining future and data covariates almost matches the full model, highlighting the significant role of the immediate temporal context in short-term predictions. ❹ Across 1-day to 3-day horizons, the most critical trend is increasing the importance of the past covariates as the forecast horizon extends, complementing the dominant predictive power of future covariates.
C.3.2 Strategy Study.
@@
Simple Weight: Each stream is processed by its own encoder 4.1. The final prediction employs fixed, equal weights:
@@
Attention: Let denote the intermediate hidden states produced by the two branches before the final read-out layer. For each directional attention we use shared projection matrices . Future Past
@@
As Figures C.1 show, we can observe that: ❶ In the 2-day horizon, ExoST maintains its advantage, while Simple Weight and Learnable Weight exhibit amplified weaknesses: the former’s static weighting fails to handle increased uncertainty in longer forecasts, and the latter, despite its adaptable weights, amplifies noise from future exogenous signals due to lacking day-to-day context modeling. ❷ Although all methods show increasing errors across horizons of 1-day to 3-day, ExoST consistently suppresses error accumulation better than alternatives.
Appendix D More Discussion
D.1 Current Limitation